A And B Are Two Vectors And Their resulants is R
  1. For What angle it is maximum
  2. for what angle is it minimumn

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
We know the expression for R=(A2+B2+2ABcos\theta)½
As A& B2 are positive and constant so the value of R will be Maximum when cos\theta is Maximum.
Hence cos\theta is maximum at \theta=0o so R is maximum at \theta=0o.
Similarly cos\theta is minimum at \theta=180o so R is minimum at \theta=180o.
If you are helped by my solution then please approve the solution to encourage me.